# Nonlinear reduction strategies for data compression: a comprehensive comparison from diffusion to advection problems

### v1

@@ 1. Introduction
In recent years, the topic of Data Reduction has become pervasive in the scientific computing community, as it contributes to popular ideas spread across many related fields, e.g. Numerical Analysis, Metric Learning, Representation Learning, Generative Modeling, and more. To pursue a comprehensive analysis of the topic, we start by defining what is meant here by reduction: identifying a low dimensional set of coordinates in the data space that encodes most of the information contained in the full dataset. The main assumption behind a reduction task is that the data manifold $\mathcal{M}$ at hand lives, with a small approximation error, on a lower dimensional manifold $\mathcal{Z}_{N}$ of
@@
Figure 2. Bi-directional flow comparing the properties gathered respectively with methods of Manifold Learning, Representation Learning, and Numerical Analysis in the field of nonlinear reduction.
In this work, we address some well-known nonlinear reduction methodologies developed in the context of Numerical Analysis, Metric Learning, and Representation Learning. Each one of them has some traditional strategies to perform data reduction, therefore, to give a glance at their connection and strengths, we report in Figure 2 a general scheme representing their links. Precisely, a double flow is proposed to report the properties gained in the reduction process passing from one methodology to another. These comments and connections will be deeply described and clarified in the following sections, nevertheless, it is worth giving a short brief overview of their main differences. Representation Learning usually builds the reduction problem as a supervised optimization of deep networks aimed at minimizing the reconstruction error of each datum independently from its neighbors. Metric Learning learns a latent space by exploiting some data-structure awareness to properly cluster the data in the new low-dimensional latent space. Finally, Reduced Order Models (ROMs) in Numerical Analysis, generally design an interpretable reduction pipeline providing convergence estimates on its reduced representations. On the contrary, learning-based algorithms usually lack interpretability, not clarifying the meaning of the obtained latent spaces, i.e. not providing information on the chosen bases for $\mathcal{Z}_{N}$ corresponding to the low-dimensional coefficients provided as latent representations of the data.
@@ 2. Challenges in linear reduction
Given the parametric data manifold :=$\mathcal{M}:=\{u_{\mu_{1}},\dots,u_{\mu_{n}}\}$, where $u_{\mu_{i}}\in\mathcal{X},\ \forall\mu_{i}\in\mathcal{P}:=\{\mu_{1},\dots,\mu_{n}\}$ with $\mathcal{X}$ Hilbert space, traditional approaches to reduction rely on linear approximation spaces of the form $\mathcal{Z}_{N}=\text{span}\{\psi_{j}\}_{j=1}^{N}$ with $\psi_{j}\in\mathcal{X},\ \forall j\in\{1,\dots,N\}$, so that
$u_{\mu_{i}}\simeq\Pi_{\mathcal{Z}_{N}}u_{\mu_{i}}:=\sum_{j=1}^{N}z_{j}(\mu_{i})\psi_{j}\quad\forall\mu_{i}\in\mathcal{P},$
given the parameter-dependent reduced variables $z_{j}:\mathcal{P}\rightarrow\mathbb{R},\ \forall j\in\{1,\dots,N\}$.
@@
While under certain circumstances, depending on the properties of $\mathcal{M}$, it is possible to demonstrate that $d_{N}(\mathcal{M})$ rapidly decays increasing $N$ [2, 4], there exists a large class of problems for which linear approximation does not offer an accurate reduced representation for a reasonably small amount “modes”. It is, for instance, the case of data collected from Advection-dominated dynamics [38, 25], which are commonly described as the prototypical linearly nonreducible $\mathcal{M}$, i.e. characterized by a slow decay of $d_{N}(\mathcal{M})$.
This can be seen by comparing the decay of PCA eigenvalues of the manifold $\mathcal{M}$ obtained by solving the following Advection-Diffusion system for different values of the parameters $c_{T}$ and $c_{D}$ in the following equation in sthe spatial domain $\Omega\subset\mathbb{R}^{D}$:
(2) $\begin{cases}\partial_{t}u(\bm{x},t)+c_{T}\nabla_{\bm{x}}\cdot u(\bm{x},t)-c_{D}\Delta_{\bm{x}}u(\bm{x},t)=0&\quad\bm{x}\in\Omega,\ t>0\\ u(\bm{x},0)=u_{0}&\quad\bm{x}\in\Omega.\end{cases}$
Indeed, for $c_{T}=0,c_{D}\neq 0$ one obtains a purely diffusive system, while for $c_{T}\neq 0,c_{D}=0$ a purly advective one. In Figure 3 we compare the eigenvalues decay of the Advection, Diffusion, and Advection-Diffusion manifolds, both obtained by collecting solutions of Equation (2) for $t\in\mathcal{P}=[0,T]$ with $T=0.5$ for fixed values of the parameters: respectively $(c_{T}=4,c_{D}=0)$, $(c_{T}=0,c_{D}=0.1)$, and $(c_{T}=4,c_{D}=0.1)$. Thus, it is clear that, even in the non-parametric setting (only evolving the systems in time), linear reduction is not well-suited for Advection-dominated problems, since the correspondent PCA eigenvalues exhibit a significantly slower decay w.r.t. the other models, for which diffusive effects are present.
@@ 3. Nonlinear Reduced Models: different Learning Strategies, same Reduction Pipeline
In this section, we address well-known nonlinear reduction strategies from different research fields by linking them to the scheme reported in Figure 1. In particular, we identify a nonlinear reduction process with a two-stage pipeline composed of a possible pre-processing step transforming the manifold $\mathcal{M}$ into $\mathcal{T}$ through the map $\Phi$, and of a standard reduction step, which aims to find the function $Z_{N}$ such that the reduced manifold $\mathcal{Z}_{N}:=Z_{N}(\mathcal{T})$ has dimension $N$ and approximates sufficiently well $\mathcal{T}$. Thus, the main goal of the first step is to obtain a “simpler” representation of the manifold $\mathcal{T}$ which is more suitable for reduction, while the second one actually searches for the more convenient $N$-dimensional description of the resulting manifold $\mathcal{T}:=\Phi(\mathcal{M})$. In addition, since it is generally desirable to reconstruct the original data $\mathcal{M}$ from $\mathcal{Z}$, the maps $\Phi$ and $Z_{N}$ should be invertible.
@@
Recalling Section 2, in the linear PCA/POD approach one can identify $\Phi$ with the identity map, so that $\mathcal{T}=\mathcal{M}$, and $Z_{N}:\Phi(u_{\mu_{i}})\mapsto\{z_{j}(\mu_{i})\}_{j=1}^{N}\in\mathcal{Z}_{N}$ with the map assigning to each transformed datum the correspondent reduced variables $\bm{z}(\mu_{i}):=\{z_{j}(\mu_{i})\}_{j=1}^{N}$ for a given parameter value $\mu_{i}\in\mathcal{P}$. Then, $Z_{N}^{-1}(\bm{z})=\sum_{i=j}^{N}z_{j}\psi_{j}$, where $\{\psi_{j}\}_{j=1}^{N}$ are defined as in (1) and correspond to the chosen bases of $\mathcal{T}(\equiv\mathcal{M})$ to project back the reduced space $\mathcal{Z}_{N}$.
@@
Now, we continue by discussing how different nonlinear reduction techniques fit in the above pipeline. With this aim, we present below the characterization of these two stages for some numerical nonlinear ROMs, for well-known Metric/Manifold Learning techniques, and for Auto-Encoders, belonging to the Representation Learning field.
3.1. Nonlinear ROMs in Numerical Analysis
@@
Nonlinear approaches to reduction, as underlined in [37, 40] can be subdivided into Eulerian and Lagrangian ones. Eulerian methods search for possibly nonlinear parameter-dependent maps $Z_{N}\equiv Z_{N,\mu}$ and $Z_{N}^{-1}\equiv Z_{N,\mu}^{-1}$, with $\mu\in\mathcal{P}$, while keeping $\Phi\equiv Id$. Lagrangian approaches instead aim to perform classical linear reduction on $\mathcal{T}=\Phi(\mathcal{M})$ with a general nonlinear map $\Phi$, obtained by properly optimizing the image set $\mathcal{T}$ to make it well-suited for linear reduction [25].
@@
Indeed, Registration Methods deal with the task of finding a parameter-dependent bijection $\Phi\equiv\Phi_{\mu}$ such that, given the initial manifold $\mathcal{M}$ of sampled solutions of a pPDE $\{u_{\mu_{i}}(\bm{x})\}_{i=1}^{n}$ with $\bm{x}\in\Omega\subset\mathbb{R}^{D}$, for an ensemble of different values of the parameter $\{\mu_{i}\}_{i=1}^{n}$, the image manifold $\mathcal{T}:=\Phi(\mathcal{M})$ is suitable for linear reduction, i.e. has a nice Kolmogorov decay. This is done by defining the optimal map $\Phi_{\mu}$ as the one that minimizes the $L_{2}$ distance between the transformed data and the reference solution $u_{\overline{\mu}}\in\mathcal{M}$, for which $\Phi_{\overline{\mu}}$ is the identity map. In particular, $\Phi_{\mu}$ is searched as an affine map of the form
@@
Thus, Registration Methods offer a general and interpretable framework, clarifying the optimization process that leads to a given choice of $\Phi$ while ensuring its invertibility and the possibility of reconstructing the original samples. Furthermore, since they rely on classical linear methods for the reduction step, they can be coupled with any reduction strategy by selecting the dimension $N$ looking at the eigenvalues decay of the covariance operator computed on the transformed manifold $\mathcal{T}=\Phi_{\mu}(\mathcal{M})$, not imposing it a-priori.
Moreover, an interesting direction for Registration has been offered by Optimal Transport (OT) theory, which focuses on the existence and computability of optimal transport maps between two probability measures with finite second order moments $\rho,\sigma\in\mathcal{P}_{2}(\Omega,\mathcal{W}_{2})$ with respect to the Wasserstein metric $\mathcal{W}_{2}$. It can be reinterpreted from a Registration perspective by considering $u_{\mu_{i}}(\bm{x})\in\mathcal{P}_{2}(\Omega,\mathcal{W}_{2}),\ \forall\mu_{i}\in\mathcal{P}$, computing $\Phi_{\mu}$ through the transport maps between the distributions $u_{\mu}(\bm{x})$ and the reference element $u_{\overline{\mu}}(\bm{x})$. Some works developed and discussed this intuition, proposing frameworks in which OT maps can be useful to find a proper $\mathcal{T}$ in case of problems with slow decay, such as the Advection-dominated ones [3, 8].
@@
[…]
3.2. Spectral Methods for Manifold Learning
Metric Learning, by definition, aims at optimizing a problem-dependent metric that provides a well-suited similarity measure to represent the given dataset by means of a few clusters. This task can be approached both in the supervised and unsupervised context, and in the latter case it also takes the name of Manifold Learning. Manifold Learning techniques aim to understand the manifold geometrical and topological features. Precisely, this is often done by identifying the bases spanning an informative lower dimensional space, on which to project the original data manifold to both preserve and emphasize possibly hidden core information. As a result, manifold learning often provides significant hints for the reduction, as it searches for a low dimensional representation of the high dimensional data while preserving relevant information of its original topology [21].
For instance, PCA (see Section 2) belongs to this class of methods, as it finds new bases maximizing the variance of the reduced representation so that the original data can be treated with a reduced number of coordinates while still preserving its intrinsic variance. A nonlinear generalization to PCA is offered by its “kernelized” version [35], which only modifies the $\Phi$. Indeed, in kernel PCA (kPCA), $\Phi$ is assumed to be a high-dimensional function defining the correspondent kernel $k$ used in computations. The sense behind a high-dimensional $\mathcal{T}$ is to obtain a sparsification of data on new directions advantageous to the subsequent PCA reduction.

### v2

@@ 1. Introduction
In recent years, the topic of Data Reduction has become pervasive in the scientific computing community, as it contributes to popular ideas spread across many related fields, e.g. Numerical Analysis, Manifold Learning, Representation Learning, Generative Modeling, and more. To pursue a comprehensive analysis of the topic, we start by defining what is meant here by reduction: identifying a low dimensional set of coordinates in the data space that encodes most of the information contained in the full dataset. The main assumption behind a reduction task is that the data manifold $\mathcal{M}$ at hand lives, with a small approximation error, on a lower dimensional manifold $\mathcal{Z}_{N}$ of
@@
Figure 2. Bi-directional flow comparing the properties gathered respectively with methods of Manifold Learning, Representation Learning, and Model Order Reduction in the context of nonlinear strategies for data compression.
In this work, we address some well-known nonlinear reduction methodologies developed in the context of Numerical Analysis with Model Order Reduction techniques, Manifold Learning, and Representation Learning. Each one of them has some traditional strategies to perform data reduction, therefore, to give a glance at their connection and strengths, we report in Figure 2 a general scheme representing their links. Precisely, a double flow is proposed to report the properties gained in the reduction process passing from one methodology to another. These comments and connections will be deeply described and clarified in the following sections, nevertheless, it is worth giving a short brief overview of their main differences. Representation Learning usually builds the reduction problem as a supervised optimization of deep networks having strong approximation capabilities demonstrated in the Universal Approximation Theorem (UAT) [25], aiming at minimizing the reconstruction error of each datum independently from its neighbors. Manifold Learning learns a latent space by exploiting some data-structure awareness to properly cluster the data in the new low-dimensional latent space. Finally, Reduced Order Models (ROMs) in Numerical Analysis, generally design an interpretable reduction pipeline providing convergence estimates on its reduced representations. On the contrary, learning-based algorithms usually lack interpretability, not clarifying the meaning of the obtained latent spaces, i.e. not providing information on the chosen bases for $\mathcal{Z}_{N}$ corresponding to the low-dimensional coefficients provided as latent representations of the data.
@@ 2. Challenges in linear reduction
Given the parametric data manifold $\mathcal{M}:=\{u_{\mu_{1}},\dots,u_{\mu_{n}}\}$, where $u_{\mu_{i}}\in\mathcal{X},\ \forall\mu_{i}\in\mathcal{P}:=\{\mu_{1},\dots,\mu_{n}\}$ with $\mathcal{X}$ Hilbert space, traditional approaches to reduction rely on linear approximation spaces of the form $\mathcal{Z}_{N}=\text{span}\{\psi_{j}\}_{j=1}^{N}$ with $\psi_{j}\in\mathcal{X},\ \forall j\in\{1,\dots,N\}$, so that
$u_{\mu_{i}}\simeq\Pi_{\mathcal{Z}_{N}}u_{\mu_{i}}:=\overline{u}+\sum_{j=1}^{N}z_{j}(\mu_{i})\psi_{j},\ \forall\mu_{i}\in\mathcal{P},$
given the parameter-dependent reduced variables $z_{j}:\mathcal{P}\rightarrow\mathbb{R},\ \forall j\in\{1,\dots,N\}$ and $\overline{u}:=\frac{1}{n}\sum_{i=1}^{n}u_{\mu_{i}}$.
@@
While under certain circumstances, depending on the properties of $\mathcal{M}$, it is possible to demonstrate that $d_{N}(\mathcal{M})$ rapidly decays for increasing $N$ [3, 8], there exists a large class of problems for which linear approximation does not offer an accurate reduced representation for a reasonably small amount “modes”. It is, for instance, the case of data collected from Advection-dominated dynamics [48, 35], which are commonly described as the prototypical linearly nonreducible $\mathcal{M}$, i.e. characterized by a slow decay of $d_{N}(\mathcal{M})$.
This can be seen by comparing the decay of PCA eigenvalues of the manifold $\mathcal{M}$ obtained by solving the following Advection-Diffusion system for different values of the parameters $c_{T}$ and $c_{D}$ in the following equation in the spatial domain $\Omega\subset\mathbb{R}^{D}$:
(2) $\begin{cases}\partial_{t}u(\bm{x},t)+c_{T}\nabla_{\bm{x}}\cdot u(\bm{x},t)-c_{D}\Delta_{\bm{x}}u(\bm{x},t)=0,&\quad\bm{x}\in\Omega,\ t>0\\ u(\bm{x},0)=u_{0},&\quad\bm{x}\in\Omega.\end{cases}$
Indeed, for $c_{T}=0,c_{D}\neq 0$ one obtains a purely diffusive system, while for $c_{T}\neq 0,c_{D}=0$ a purely advective one. Note that, when dealing with time-dependent systems, one may choose to include the time variable in the parameters’ set. In Figure 3 we compare the eigenvalues decay of the Advection, Diffusion, and Advection-Diffusion manifolds, both obtained by collecting solutions of Equation (2) for $t\in\mathcal{P}=[0,T]$ with $T=0.5$ for fixed values of the other parameters: respectively $(c_{T}=4,c_{D}=0)$, $(c_{T}=0,c_{D}=0.1)$, and $(c_{T}=4,c_{D}=0.1)$. Thus, it is clear that, even in the “non-parametric setting” (only evolving the systems in time), linear reduction is not well-suited for Advection-dominated problems, since the correspondent PCA eigenvalues exhibit a significantly slower decay w.r.t. the other models, for which diffusive effects are present.
@@ 3. Nonlinear Reduced Models: different Learning Strategies, same Reduction Pipeline
In this section, we address well-known nonlinear reduction strategies from different research fields by linking them to the scheme reported in Figure 1. In particular, we identify a nonlinear reduction process with a two-stage pipeline composed of a possible pre-processing step transforming the manifold $\mathcal{M}$ into $\mathcal{T}$ through the operator $\Phi:\mathcal{M}\rightarrow\mathcal{T}$, and of a standard reduction step, which aims to find the operator $Z_{N}:\mathcal{T}\rightarrow\mathcal{Z}_{N}\subseteq\mathbb{R}^{N}$ such that the reduced manifold $\mathcal{Z}_{N}:=Z_{N}(\mathcal{T})$ has dimension $N$ and approximates sufficiently well $\mathcal{T}$. Thus, the main goal of the first step is to obtain a “simpler” representation of the manifold $\mathcal{T}$ which is more suitable for reduction, while the second one actually searches for the more convenient $N$-dimensional description of the resulting manifold $\mathcal{T}:=\Phi(\mathcal{M})$. For instance, in the case of projection-based ROMs, the last step aims to find $N$ bases $\{\psi_{j}\}_{j=1}^{N}$ on which to project $\mathcal{T}$, so that $Z_{N}$ is taken as a projection operator. In addition, since it is generally desirable to reconstruct the original data $\mathcal{M}$ from $\mathcal{Z}_{N}$, the methods should also provide the operators $\Phi^{-1}:\mathcal{T}\rightarrow\mathcal{M}$ and $Z_{N}^{-1}:\mathcal{Z}_{N}\rightarrow\mathcal{T}$. Nevertheless, we underline the abuse of notation for $\Phi^{-1}$ and $Z_{N}^{-1}$, as they are defined only to provide a way to map back the reduced representation to the original manifold, being $\Phi$ and $Z_{N}$ not necessarily invertible operators. Therefore, each method gives its own characterization to $\Phi^{-1}$ and $Z_{N}^{-1}$ (see for instance Section 3.3 for an example of non-invertible $Z_{N}$).
@@
Recalling Section 2, in the linear PCA/POD approach one can identify $\Phi$ with the identity operator, so that $\mathcal{T}=\mathcal{M}$, and $Z_{N}:\Phi(u_{\mu_{i}})\mapsto\{z_{j}(\mu_{i})\}_{j=1}^{N}\in\mathcal{Z}_{N}$ with the operator assigning to each transformed datum the reduced variables $\bm{z}(\mu_{i}):=\{z_{j}(\mu_{i})\}_{j=1}^{N}$ for a given parameter value $\mu_{i}\in\mathcal{P}$ corresponding to the projection of $\Phi(u_{\mu_{i}})$ on the bases $\psi_{j}$. Then the “inverse” operator $Z_{N}^{-1}(\bm{z})=\sum_{i=j}^{N}z_{j}\psi_{j}$ projects back the reduced variables on $\mathcal{T}$, being $\{\psi_{j}\}_{j=1}^{N}$ defined as in (1) and corresponding to the chosen bases of $\mathcal{T}(\equiv\mathcal{M})$. Moreover, we highlight that the same reasoning applies for methods that rely on the interpolation of different local projection-based ROMs correspondent to different parameters values [1, 36]. Indeed, such approaches can be represented by taking $\Phi$ as the identity map (similarly to the classical PCA case), and choosing $Z_{N}$ to be a projection-operator on a set of parameter dependent bases $\{\psi_{j}(\mu)\}_{j=1}^{N}$, so that $\mathcal{Z}_{N}\equiv\mathcal{Z}_{N}(\mu)$.
@@
Now, we continue by discussing how different nonlinear reduction techniques fit in the above pipeline. With this aim, we present below the characterization of these two stages for some numerical nonlinear ROMs, for well-known Manifold Learning techniques, and for Auto-Encoders, belonging to the Representation Learning field.
3.1. Techniques from Model Order Reduction: Registration methods
@@
Nonlinear approaches to reduction, as underlined in [50, 54], can be subdivided into Eulerian and Lagrangian ones. Eulerian methods search for possibly nonlinear parameter-dependent maps $Z_{N}\equiv Z_{N,\mu}$ and $Z_{N}^{-1}\equiv Z_{N,\mu}^{-1}$, with $\mu\in\mathcal{P}$, while keeping $\Phi\equiv Id$. Lagrangian approaches instead aim to perform classical linear reduction on $\mathcal{T}=\Phi(\mathcal{M})$ with a general nonlinear map $\Phi$, obtained by properly optimizing the image set $\mathcal{T}$ to make it well-suited for linear reduction [35].
@@
Indeed, Registration Methods [13, 49, 51] deal with the task of finding a parameter-dependent bijection $\Phi\equiv\Phi_{\mu}$ such that, given the initial manifold $\mathcal{M}$ of sampled solutions of a pPDE $\{u_{\mu_{i}}(\bm{x})\}_{i=1}^{n}$ with $\bm{x}\in\Omega\subset\mathbb{R}^{D}$, for an ensemble of different values of the parameter $\{\mu_{i}\}_{i=1}^{n}$, the image manifold $\mathcal{T}:=\Phi(\mathcal{M})$ is suitable for linear reduction, i.e. has a nice Kolmogorov decay. This is done by defining the optimal map $\Phi_{\mu}$ as the one that minimizes the $L_{2}$ distance between the transformed data and the reference solution $u_{\overline{\mu}}\in\mathcal{M}$, for which $\Phi_{\overline{\mu}}$ is the identity map. In particular, $\Phi_{\mu}$ is searched as an affine map of the form
@@
Thus, Registration Methods offer a general and interpretable framework, clarifying the optimization process that leads to a given choice of $\Phi$ while ensuring its invertibility and the possibility of reconstructing the original samples. Furthermore, since they rely on classical linear methods for the reduction step, they can be coupled with any reduction strategy by selecting the dimension $N$ looking at the eigenvalues decay of the covariance operator computed on the transformed manifold $\mathcal{T}=\Phi(\mathcal{M}):=\{u_{\mu}\circ\Phi_{\mu}^{\hat{a}}\ |\ u_{\mu}\in\mathcal{M}\}$, not imposing it a-priori.
Moreover, an interesting direction for Registration has been offered by Optimal Transport (OT) theory, which focuses on the existence and computability of optimal transport maps between two probability measures with finite second order moments $\rho,\sigma\in\mathcal{P}_{2}(\Omega,\mathcal{W}_{2})$ with respect to the Wasserstein metric $\mathcal{W}_{2}$. It can be reinterpreted from a Registration perspective by considering $u_{\mu_{i}}(\bm{x})\in\mathcal{P}_{2}(\Omega,\mathcal{W}_{2}),\ \forall\mu_{i}\in\mathcal{P}$, computing $\Phi$ through the transport maps $T_{u_{\mu}\rightarrow u_{\overline{\mu}}}$ between the distributions $u_{\mu}(\bm{x})$ and the reference element $u_{\overline{\mu}}(\bm{x})$. Some works developed and discussed this intuition, proposing frameworks in which OT maps can be useful to find a proper $\mathcal{T}=\Phi(\mathcal{M}):=\{(T_{u_{\mu}\rightarrow u_{\overline{\mu}}})_{\#}u_{\mu}\ |\ u_{\mu}\in\mathcal{M}\}$ in case of problems with slow decay, such as the Advection-dominated ones [7, 12].
@@
$\mathcal{W}_{2}(u_{\mu},u_{\overline{\mu}})=\min_{T}\int_{\Omega}\frac{1}{2}|T(u_{\mu})-u_{\overline{\mu}}|^{2}d\bm{x}.$
3.2. Techniques from Manifold Learning: Kernel and Spectral methods
Manifold Leaning belongs to the Metric Learning field, which, by definition, aims at optimizing a problem-dependent metric that provides a well-suited similarity measure to represent the given dataset by means of a few clusters. This task can be approached both in the supervised and unsupervised context, and in the latter case it precisely takes the name of Manifold Learning. Such techniques aim to understand the manifold geometrical and topological features. Precisely, this is often done by identifying the bases spanning an informative lower dimensional space, on which to project the original data manifold to both preserve and emphasize possibly hidden core information. As a result, manifold learning often provides significant hints for the reduction, as it searches for a low dimensional representation of the high dimensional data while preserving relevant information of its original topology [30, 27, 22, 5].
For instance, PCA (see Section 2) belongs to this class of methods, as it finds new bases maximizing the variance of the reduced representation so that the original data can be treated with a reduced number of coordinates while still preserving its intrinsic variance. A nonlinear generalization to PCA is offered by its “kernelized” version [46, 6], which only modifies the $\Phi$. Indeed, in kernel PCA (kPCA), $\Phi$ is assumed to be a high-dimensional function defining the correspondent kernel $k$ used in computations. The idea behind a high-dimensional $\mathcal{T}$ is to obtain a sparsification of data on new directions advantageous to the subsequent PCA reduction.